# In celebration of Prof. Xun Yu Zhou’s 60th birthday

### v1

@@ 2 Preliminary
To fit the transfer learning into our framework, we can take ${\mu_{\rm ini}}=p_{\rm sou}$, ${\mu_{\rm tar}}=p_{\rm tar}$, and set $b\equiv 0$ for simplicity. In this case, if we choose $\alpha=0$, and $X_{0}\sim p_{\rm sou}$, then $X^{0}_{T}=X_{0}+W_{T}\sim p_{\rm sou}*\mathcal{N}(0,T\mathbb{I}_{d})$, where $\mathcal{N}(0,T{\mathbb{I}_{d}})=\mathbb{P}^{0}\circ W_{T}^{-1}$ and $\mathbb{I}_{d}$ denotes the $d\times d$ identity matrix. Thus, denoting the optimal control by $\widehat{\alpha}$ and noting that $\alpha\equiv 0$ is sub-optimal, we must have
@@ 3 Existence of Optimal Policies for SCSBP’s
In this section we study the stochastic control problem (2.9)-(2.10) and the associated soft-constrained SBP. In particular, we shall prove that the optimal control for each $k\in\mathbb{N}$ exists and in next section we will show that these optimal policies will converge to the solution of the original SBP, with a linear rate of convergence. We shall assume that the target distribution for the SCSBP has a density $f_{\rm tar}\in\mathbb{L}^{1}(\mathbb{R}^{d})$. Also, we shall assume $\sigma(\cdot)=I_{d}$, that is, in what follows we assume that the underlying diffusion takes the form:
$dX_{t}=b(t,X_{t})dt+dW_{t},\quad X_{0}\sim{\mu_{\rm ini}},\quad t\in[0,T],$ (3.1)
where $W$ is the canonical Brownian motion under $\mathbb{P}^{0}$. Let $p(\cdot,\cdot;\cdot,\cdot)$ denote the transition density of the solution $X$, so that $\mathbb{P}^{0}\{X_{s}\in dz|X_{t}=x\}=p(s,z;t,x)dz$, $0\leq t<s\leq T$, $z,x\in\mathbb{R}^{d}$. Then, it is well known that $p(\cdot,\cdot;\cdot,\cdot)$ is the fundamental solution to Kolmogorov backward (parabolic) PDE, and under mild conditions (see, e.g., [3]), there exist $c_{1}$, $c_{2}$, $\lambda$, $\Lambda>0$, it holds that
@@
Keeping the original SBP (2.4) and (2.5) associated with (3.1) in mind, let us now recall the Problem 2.4 and the cost functional $J^{k}(\alpha)$ defined by (2.10). For notational simplicity in what follows we shall identify $\mathbb{Q}\in{\cal A}$ by $\mathbb{Q}=\mathbb{P}_{X^{\alpha}}=\mathbb{P}^{0}\circ(X^{\alpha})^{-1}$. Clearly, we have $\mathbb{P}_{X^{0}}=\mathbb{P}_{X}$, where $X$ solves (3.1). Furthermore, we shall denote $\mathbb{E}[\cdot]=\mathbb{E}^{\mathbb{P}^{0}}[\cdot]$ when context is clear, and for each $k\in\mathbb{N}$ we can easily check that
@@
$h(t,x)=\int_{\mathbb{R}^{d}}p(T,z;t,x)\frac{f_{\rm tar}(z)}{p(T,z;0,x_{0})}dz:=\mathbb{E}\Big[\frac{f_{\rm tar}(X_{T})}{p(T,X_{T};0,x_{0})}\Big|X_{t}=x\Big],$ (3.5)
@@
Another natural example of $G$ satisfying Assumption 3.2 would be the Wasserstein distance or the KL divergence, augmented with a small “guardrail” term that enforces the uniform (weighted) pointwise control in (3.7). This guardrail can be taken as a weighted $L_{\infty}$ norm, a Hölder $\mathbb{C}^{\alpha}$ seminorm, or an RKHS norm (e.g., with kernel $k(x,y)=\phi(x)\phi(y)\kappa(x-y)$). In next example we illustrate such a choice with the Wasserstein distance plus an $L_{\infty}$ guardrail.
@@
$|f_{\mu}(x)-f_{\rm tar}(x)|\leq\Big\|\frac{f_{\mu}-f_{\rm tar}}{\phi}\Big\|_{L^{\infty}}\phi(x)\leq\frac{1}{c}\phi(x)G(\mu).$
@@
Let $\{\mu_{n}\}_{n\geq 1}\subset\mathscr{P}_{2}(\mathbb{R}^{d})$ with $\|\mu_{n}\|^{2+\varepsilon}\rightarrow\infty$. We claim that $G(\mu_{n})$ must be unbounded. Indeed, suppose not. Then there exists $M,M^{\prime}>0$ such that $W_{2}^{2}(\mu_{n},{\mu_{\rm tar}})\leq M$ and $\|\phi^{-1}(f_{\mu_{n}}-f_{\rm tar})\|_{L^{\infty}}<M^{\prime}$ for all $n\in\mathbb{N}_{+}$. Hence $f_{\mu_{n}}(x)\leq f_{\rm tar}(x)+M^{\prime}\phi(x)$, $x\in\mathbb{R}^{d}$. Integrating against $|x|^{2+\varepsilon}$ and using the facts that ${\mu_{\rm tar}}\in\mathscr{P}_{2+\varepsilon}(\mathbb{R}^{d})$ with $\varepsilon=p-2>0$ and $\int|x|^{2+\varepsilon}\phi(x)dx<\infty$, we have
@@
Since $\inf_{\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})}D_{k}(\mu)\leq D_{k}({\mu_{\rm tar}})=m$, we may assume without loss of generality that
$\{\mu_{k}^{(i)}\}_{i=1}^{\infty}\subset\mathcal{S}_{k}$. Since $\mathcal{S}_{k}$ is uniformly integrable and is tight in $\mathscr{P}_{2}(\mathbb{R}^{d})$, there exists subsequence $\{\mu_{k}^{(i_{l})}\}_{l=1}^{\infty}$ such that $\widehat{\mu}_{k}:=\lim_{l\rightarrow\infty}\mu_{k}^{(i_{l})}\in\mathscr{P}_{2}(\mathbb{R}^{d})$ [^3] This follows from the result on Wasserstein distance vs. weak convergence (see, e.g., [65, Theorem 7.12]), which states that $W_{p}(\mu_{k},\mu)\to 0$ if and only if $\mu_{k}\to\mu$ weakly, and $\lim_{R\to\infty}\mathop{\overline{\rm lim}}_{k\to\infty}\int_{\{d(x,x_{0})\geq R\}}d(x,x_{0})^{p}\mu_{k}(dx)=0$.. Since the mapping $\mu\mapsto D_{k}(\mu)$ is continuous, we have
@@
proving the claim. Furthermore, if we denote the density of $\mathbb{P}_{X_{T}}$ by $f_{X_{T}}$, and note that $D_{\rm KL}(\widehat{\mu}_{k};\mathbb{P}_{X_{T}})\leq D_{k}(\widehat{\mu}_{k})\leq m<\infty$, we know that $\frac{d\widehat{\mu}_{k}}{d\mathbb{P}_{X_{T}}}$ exists and
@@
Let $k\in\mathbb{N}$ be fixed, and let $\widehat{\mu}_{k}$ be the minimizer of $D_{k}(\cdot)$ defined by (3.8). Then, by (3.4), for any $\alpha\in\mathbb{L}^{2}_{\mathbb{F}^{0}}([0,T])$, we have
@@ 4 Convergence Results under Delta Initial Distribution
We make the following two observations. First, if we denote $g(x):=\frac{f_{\rm tar}(x)}{p(T,x;0,x_{0})}$, then by (3.5) we can write $h(t,x)=\mathbb{E}_{t,x}[g(X_{T})]:=\mathbb{E}[g(X_{T})|X_{t}=x]$, where $X$ is the solution to (3.1) with $X_{0}=x_{0}$. By Feynman-Kac formula, we see that $h$ satisfy the PDE:
@@
Recall that $\widehat{\mu}_{k}\in{\cal S}_{k}$, we have $kG(\widehat{\mu}_{k})\leq D_{k}(\widehat{\mu}_{k})\leq m$, or
@@
Then Assumption 3.2-(ii) amounts to saying that $|f_{k}(x)-f_{\rm tar}(x)|\leq\frac{C^{\prime}}{k}$ with constant $C^{\prime}=CmM$. In other words, for all $x\in\mathbb{R}^{d}$, as $k\to\infty$ we have
$\left\{\begin{array}[]{lll}\displaystyle|g(x)-g_{k}(x)|=|h(T,x)-h^{k}(T,x)|=\frac{1}{p(T,x;0,x_{0})}|f_{\rm tar}(x)-f_{k}(x)|\to 0\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \displaystyle|h(t,x)-h^{k}(t,x)|\leq\int_{\mathbb{R}^{d}}\frac{p(T,z;t,x)}{p(T,z;0,x_{0})}|f_{\rm tar}(z)-f_{k}(z)|dz\to 0,\end{array}\right$.
We shall use these facts to study the convergence of the optimal policies in the next subsection.
@@ 4.1 The Convergence of Optimal Policies
$\leq$ $\left|\frac{\nabla h^{k}(t,x)-\nabla h(t,x)}{h^{k}(t,x)}\right|+|\nabla h(t,x)|\left|\frac{h(t,x)-h^{k}(t,x)}{h^{k}(t,x)h(t,x)}\right|:=I_{1}+I_{2}.$
@@
$h^{k}(t,x)=d_{k}^{-1}\int p(T,z,;t,x)\Big(\frac{f_{\rm tar}(z)}{p(T,z;0,x_{0})}\Big)^{\frac{k}{k+1}}dz,$
with $d_{k}=\int f_{\rm tar}(x)^{\frac{k}{1+k}}p(x,T\,|\,x_{0},0)^{\frac{1}{1+k}}dx$. In addition, $J^{k}(\widehat{\alpha})=-(1+k)\,\log(C_{k})$. Consequently, Assumption 3.2-(ii) can be reduced to that $\mathbb{E}\Big[\big|\frac{f_{\rm tar}(X_{T})}{p(T,X_{T};0,x_{0})}\big|^{2}\Big]$ is bounded (see Assumption 5.4 for similar conditions); and the linear rate of convergence can be proved with the same arguments. ∎
@@ 4.2 The Convergence of the Value Function
$\negthinspace\leq\negthinspace$ $\negthinspace\frac{c}{k}\mathbb{E}\Big[\frac{1}{2}\int_{0}^{T-\varepsilon}\frac{1}{\sqrt{T-s}}\big(|\widehat{\alpha}^{k}_{s}|+|\widehat{\alpha}_{s}|\big)ds\Big]\leq\frac{c}{k\sqrt{\varepsilon}}\mathbb{E}\Big[\frac{1}{2}\int_{0}^{T}\big(|\widehat{\alpha}^{k}_{s}|+|\widehat{\alpha}_{s}|\big)\,ds\Big],$
@@
Thus, since $\widehat{\alpha}$ satisfies the constraint dynamics (2.6), we have
@@
(i) The linear (i.e., $\sim\frac{1}{k}$) "closeness" between the law of the optimal state and $\mu_{\rm tar}$} has appeared several times so far. For example, (4.3) implies that $G(\mathbb{P}_{X^{\widehat{\alpha}^{k}}_{T}})=G(\mathbb{P}_{X^{\widehat{\alpha}^{k}}_{T}};{\mu_{\rm tar}})\leq\frac{c}{k}$, and by Remark 3.3-(ii), this implies that $W_{1}(\mathbb{P}_{X^{\widehat{\alpha}^{k}}_{T}},{\mu_{\rm tar}})\sim\frac{1}{k}$. The result in (4.19) is in the same spirit, by under the stronger ${W}_{2}$-distance, but compensated by an early stopping.
@@
$\begin{cases}dX^{\widehat{\alpha}}_{t}=[b(t,X^{\widehat{\alpha}}_{t})+\widehat{\alpha}_{t}(X^{\widehat{\alpha}}_{t})]dt+dW_{t}\qquad\quad&X^{\widehat{\alpha}}_{0}=x_{0};\\ dX^{\widehat{\alpha}^{k}}_{t}=[b(t,X^{\widehat{\alpha}^{k}}_{t})+\widehat{\alpha}_{t}^{k}(X^{\widehat{\alpha}^{k}}_{t})]dt+dW_{t},&X^{\widehat{\alpha}}_{0}=x_{0}.\end{cases}$ (4.20)
@@
$\begin{cases}dX^{\widehat{\alpha}}_{t}=b^{\widehat{\alpha}}(t,X^{\widehat{\alpha}}_{t})dt+dW_{t}\qquad\quad&X^{\widehat{\alpha}}_{0}=x_{0};\\ dX^{\widehat{\alpha}^{k}}_{t}=[b^{\widehat{\alpha}}(t,X^{\widehat{\alpha}^{k}}_{t})+\Delta\widehat{\alpha}_{t}^{k}(X^{\widehat{\alpha}^{k}}_{t})]dt+dW_{t},&X^{\widehat{\alpha}}_{0}=x_{0}.\end{cases}$

### v2

@@ 2 Preliminary
To fit the transfer learning into our framework, we can take ${\mu_{\rm ini}}=p_{\rm sou}$, ${\mu_{\rm tar}}=p_{\rm tar}$, and set $b\equiv 0$ for simplicity. In this case, if we choose $\alpha=0$, and $X_{0}\sim p_{\rm sou}$, then $X_{T}=X_{0}+W_{T}\sim p_{\rm sou}*\mathcal{N}(0,T\mathbb{I}_{d})$, where $\mathcal{N}(0,T{\mathbb{I}_{d}})=\mathbb{P}^{0}_{W_{T}}$ and $\mathbb{I}_{d}$ denotes the $d\times d$ identity matrix. Thus, denoting the optimal control by $\widehat{\alpha}$ and noting that $\alpha\equiv 0$ is sub-optimal, we must have
@@ 3 Existence of Optimal Policies for SCSBP’s
In this section we study the stochastic control problem (2.9)-(2.10) and the associated soft-constrained SBP. In particular, we shall prove that the optimal control for each $k\in\mathbb{N}$ exists and in next section we will show that these optimal policies will converge to the solution of the original SBP, with a linear rate of convergence. We shall assume that the target distribution for the SCSBP has density $f_{\rm tar}\in\mathbb{L}^{1}(\mathbb{R}^{d})$. Also, we assume $\sigma(\cdot)=I_{d}$, that is, we consider the following SDE:
$dX_{t}=b(t,X_{t})dt+dW_{t},\quad{\mathbb{P}}_{X_{0}}={\mu_{\rm ini}}.$ (3.1)
In what follows we shall often consider the SDE with initial data $(t,x)\in[0,T]\times\mathbb{R}^{d}$:
$X_{s}=x+\int_{t}^{s}b(r,X_{r})dr+W_{s}-W_{t},\quad s\in[t,T].$ (3.2)
We denote the solution of (3.2) by $X^{t,x}$. Clearly, if we let $p(\cdot,\cdot;\cdot,\cdot)$ be the transition density of the solution $X$ to (3.1), that is, $\mathbb{P}\{X_{s}\in dz|X_{t}=x\}=p(s,z;t,x)dz$, $0\leq t<s\leq T$, $z,x\in\mathbb{R}^{d}$, then $p(s,\cdot;t,x)$ is the density of $X^{t,x}_{s}$. It is well known that $p(\cdot,\cdot;\cdot,\cdot)$ is the fundamental solution to Kolmogorov backward (parabolic) PDE, and under mild conditions (see, e.g., [3]), there exist $c_{1}$, $c_{2}$, $\lambda$, $\Lambda>0$, it holds that
@@
Keeping the original SBP (2.5) associated with (3.1) in mind, let us now recall the Problem 2.4 and the cost functional $J^{k}(\alpha)$ defined by (2.10). For notational simplicity, we let $\mathbb{P}_{X^{\alpha}}$ (resp. $\mathbb{P}_{X}$) denote the probability measure on $\mathbb{C}^{d}_{T}$ induced by $X^{\alpha}$ (resp. $X$). Clearly, when $\alpha\equiv 0$, we have $\mathbb{P}_{X^{0}}=\mathbb{P}_{X}$, where $X$ solves (3.1). Furthermore, we shall denote $\mathbb{E}[\cdot]=\mathbb{E}^{\mathbb{P}}[\cdot]$ when context is clear, and for each $k\in\mathbb{N}$ we can easily check that
@@
$h(t,x)=\int_{\mathbb{R}^{d}}p(T,z;t,x)\frac{f_{\rm tar}(z)}{p(T,z;0,x_{0})}dz=\mathbb{E}\Big[\frac{f_{\rm tar}(X_{T})}{p(T,X_{T};0,x_{0})}\Big|X_{t}=x\Big],$ (3.6)
@@
Another natural example of $G$ satisfying Assumption 3.2 would be the Wasserstein distance or the KL divergence, augmented with a small “guardrail” term that enforces the uniform (weighted) pointwise control in (3.8). This guardrail can be taken as a weighted $\mathbb{L}_{\infty}$ norm, a Hölder $\mathbb{C}^{\alpha}$ seminorm, or an RKHS norm (e.g., with kernel $k(x,y)=\phi(x)\phi(y)\kappa(x-y)$). In next example we illustrate such a choice with the Wasserstein distance plus an $\mathbb{L}_{\infty}$ guardrail.
@@
$|f_{\mu}(x)-f_{\rm tar}(x)|\leq\Big\|\frac{f_{\mu}-f_{\rm tar}}{\phi}\Big\|_{\mathbb{L}^{\infty}}\phi(x)\leq\frac{1}{c}\phi(x)G(\mu).$
@@
Let $\{\mu_{n}\}_{n\geq 1}\subset\mathscr{P}_{2}(\mathbb{R}^{d})$ with $\|\mu_{n}\|^{2+\varepsilon}\rightarrow\infty$. We claim that $G(\mu_{n})$ must be unbounded. Indeed, suppose not. Then there exists $M,M^{\prime}>0$ such that $W_{2}^{2}(\mu_{n},{\mu_{\rm tar}})\leq M$ and $\|\phi^{-1}(f_{\mu_{n}}-f_{\rm tar})\|_{\mathbb{L}^{\infty}}<M^{\prime}$ for all $n\in\mathbb{N}_{+}$. Hence $f_{\mu_{n}}(x)\leq f_{\rm tar}(x)+M^{\prime}\phi(x)$, $x\in\mathbb{R}^{d}$. Integrating against $|x|^{2+\varepsilon}$ and using the facts that ${\mu_{\rm tar}}\in\mathscr{P}_{2+\varepsilon}(\mathbb{R}^{d})$ with $\varepsilon=p-2>0$ and $\int|x|^{2+\varepsilon}\phi(x)dx<\infty$, we have
@@
Since $\inf_{\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})}D_{k}(\mu)\leq D_{k}({\mu_{\rm tar}})=m$, we may assume without loss of generality that $\{\mu_{k}^{(i)}\}_{i=1}^{\infty}\subset\mathcal{S}_{k}$. Since $\mathcal{S}_{k}$ is uniformly integrable and is tight in $\mathscr{P}_{2}(\mathbb{R}^{d})$, there exists subsequence $\{\mu_{k}^{(i_{l})}\}_{l=1}^{\infty}$ such that $\widehat{\mu}_{k}:=\lim_{l\rightarrow\infty}\mu_{k}^{(i_{l})}\in\mathscr{P}_{2}(\mathbb{R}^{d})$.  [^3] This follows from the result on Wasserstein distance vs. weak convergence (see, e.g., [65, Theorem 7.12]), which states that $W_{p}(\mu_{k},\mu)\to 0$ if and only if $\mu_{k}\Rightarrow\mu$ weakly, and $\lim_{R\to\infty}\mathop{\overline{\rm lim}}_{k\to\infty}\int_{\{|x-x_{0}|\geq R\}}|x-x_{0}|^{p}\mu_{k}(dx)=0$.. Since the mapping $\mu\mapsto D_{k}(\mu)$ is continuous, we have
@@
proving the claim. Furthermore, if we denote the density of $\mathbb{P}_{X_{T}}$ by $f_{X_{T}}$, and note that $D_{\rm KL}(\widehat{\mu}_{k}\|\mathbb{P}_{X_{T}})\leq D_{k}(\widehat{\mu}_{k})\leq m<\infty$, we know that $\frac{d\widehat{\mu}_{k}}{d\mathbb{P}_{X_{T}}}$ exists and
@@
Let $k\in\mathbb{N}$ be fixed, and let $\widehat{\mu}_{k}$ be the minimizer of $D_{k}(\cdot)$ defined by (3.9). Then, by (3.5), for any $\alpha\in\mathbb{L}^{2}_{\mathbb{F}^{0}}([0,T];\mathbb{R}^{d})$, we have
@@ 4 Convergence Results under Delta Initial Distribution
To show the convergence of both policies and value functions, we first make the following observations. First, if we denote $g(x):=\frac{f_{\rm tar}(x)}{p(T,x;0,x_{0})}$ and $\mathbb{E}_{t,x}[\cdot]:=\mathbb{E}[\cdot|X_{t}=x]$, then by (3.6) we can write $h(t,x)=\mathbb{E}_{t,x}[g(X_{T})]$, where $X$ is the solution to (3.1) with $X_{0}=x_{0}$. By Feynman-Kac formula, we see that $h$ satisfy the PDE:
@@
Another useful observation is that for $\widehat{\mu}_{k}\in{\cal S}_{k}$, we have $kG(\widehat{\mu}_{k})\leq D_{k}(\widehat{\mu}_{k})\leq m$, or
@@
With the above observations, the key is to establish the convergence rate of $|h(t,x)-h^{k}(t,x)|$ and $|\nabla h(t,x)-\nabla h^{k}(t,x)|$ as $k$ tends to infinity.
@@ 4.1 The Convergence of Optimal Policies
$\leq$ $\left|\frac{\nabla h^{k}(t,x)-\nabla h(t,x)}{h^{k}(t,x)}\right|+|\nabla h(t,x)|\left|\frac{h(t,x)-h^{k}(t,x)}{h^{k}(t,x)h(t,x)}\right|=:I_{1}+I_{2}.$
@@
$h^{k}(t,x)=d_{k}^{-1}\int p(T,z;t,x)\Big(\frac{f_{\rm tar}(z)}{p(T,z;0,x_{0})}\Big)^{\frac{k}{k+1}}dz,$
with $d_{k}=\int f_{\rm tar}(x)^{\frac{k}{1+k}}p(x,T;\,x_{0},0)^{\frac{1}{1+k}}dx$. Consequently, Assumption 3.2-(ii) can be reduced to that $\mathbb{E}\Big[\big|\frac{f_{\rm tar}(X_{T})}{p(T,X_{T};0,x_{0})}\big|^{2}\Big]$ is bounded (see Assumption 5.4 for similar conditions); and the linear rate of convergence can be proved with the same arguments. ∎
@@ 4.2 The Convergence of the Value Function
$\negthinspace\leq\negthinspace$ $\negthinspace{\frac{C}{k}}\mathbb{E}\Big[\frac{1}{2}\int_{0}^{T-\varepsilon}\frac{1}{\sqrt{T-s}}\big(|\widehat{\alpha}^{k}_{s}|+|\widehat{\alpha}_{s}|\big)ds\Big]$
$\leq$ $\frac{C}{k\sqrt{\varepsilon}}\mathbb{E}\Big[\frac{1}{2}\int_{0}^{T}\big(|\widehat{\alpha}^{k}_{s}|+|\widehat{\alpha}_{s}|\big)\,ds\Big],$
@@
Thus, since $\widehat{\alpha}$ satisfies the constrained dynamics (2.6), we have
@@
(i) The linear (i.e., $\sim\frac{1}{k}$) “closeness” between the law of the optimal state and $\mu_{\rm tar}$} has appeared several times so far. For example, (4.3) implies that $G(\mathbb{P}_{X^{\widehat{\alpha}^{k}}_{T}})=G(\mathbb{P}_{X^{\widehat{\alpha}^{k}}_{T}};{\mu_{\rm tar}})\leq\frac{c}{k}$, and by Remark 3.3-(ii), this implies that $W_{1}(\mathbb{P}_{X^{\widehat{\alpha}^{k}}_{T}},{\mu_{\rm tar}})\sim\frac{1}{k}$. The result in (4.18) is in the same spirit, by under the stronger ${W}_{2}$-distance, but compensated by an early stopping.
@@
$\begin{cases}dX^{\widehat{\alpha}}_{t}=[b(t,X^{\widehat{\alpha}}_{t})+\widehat{\alpha}_{t}(X^{\widehat{\alpha}}_{t})]dt+dW_{t}\qquad\quad&X^{\widehat{\alpha}}_{0}=x_{0};\\ dX^{\widehat{\alpha}^{k}}_{t}=[b(t,X^{\widehat{\alpha}^{k}}_{t})+\widehat{\alpha}_{t}^{k}(X^{\widehat{\alpha}^{k}}_{t})]dt+dW_{t},&X^{\widehat{\alpha}^{k}}_{0}=x_{0}.\end{cases}$ (4.19)
@@
$\begin{cases}dX^{\widehat{\alpha}}_{t}=b^{\widehat{\alpha}}(t,X^{\widehat{\alpha}}_{t})dt+dW_{t}\qquad\quad&X^{\widehat{\alpha}}_{0}=x_{0};\\ dX^{\widehat{\alpha}^{k}}_{t}=[b^{\widehat{\alpha}}(t,X^{\widehat{\alpha}^{k}}_{t})+\Delta\widehat{\alpha}_{t}^{k}(X^{\widehat{\alpha}^{k}}_{t})]dt+dW_{t},&X^{\widehat{\alpha}^{k}}_{0}=x_{0}.\end{cases}$